OpenStudy (destinymasha):

between which two consecutive integers does the square root of 155 lie?

OpenStudy (zehanz):

Try to look for "easy" squares: 10²=100 11²=121 12²=... 13²=... 14²=... You will see that 155 is somewhere between these numbers. Then you also will know the answer...

OpenStudy (destinymasha):

???

OpenStudy (destinymasha):

i just didn't understand what exactly a consecutive integer.

OpenStudy (zehanz):

Consecutive integers are integers that differ 1, so 17 and 18, 11 and 12 etc. Let's call \(a=\sqrt{155}\), then \(a^2=155\), so we are looking for a number that, when squared, gives 155. So I tried to square some numbers, to see if I could get near 155.

OpenStudy (destinymasha):

ok

OpenStudy (destinymasha):

but the square root is 12.449899598?

OpenStudy (destinymasha):

do i round?

OpenStudy (zehanz):

This is not about using a calculator. You don't need it at all. You are only asked to give a rough estimation of \(\sqrt{155}\), by finding two consecutive integers. As 12²=144 (too small), and 13²=169 (too big), the answer you have to give is : "between 12 and 13". There is no rounding off. You can do this in your head, or with a piece of paper and a pencil.

OpenStudy (destinymasha):

Ooohh! Ok! Seems like i was making it harder than it already was. Thanks!

OpenStudy (zehanz):

YW!